# Individual causal effect estimation accounting for latent disease state modification among bipolar participants in mobile health studies

### v1

@@ 4 Model implementation
We consider a frequentist hidden Markov model approach (HMM-F) to estimate the parameters defined in Section 3. We note in Appendix A we explore an equivalent model fit within a Bayesian framework using a Markov Chain Monte Carlo sampling procedure. For the frequentist model, we implement an expectation-maximization (EM) procedure based on the Baum-Welch algorithm. In the expectation step, forward probabilities $\alpha_{i}(t)$ and backward probabilities $\beta_{i}(t)$ are calculated recursively using $a_{ij,t-1}$ and $b_{j,Y_{t},A_{t},C_{t}}(y_{t},a_{t},c_{t})$ [46, 28] such that
@@
To update the $\beta$, $\psi$, and $\phi$ estimates, we use the appropriate generalized linear regression for $Y_{t}$, $A_{t}$, and $C_{t}$ as the outcome with $\gamma_{i}(t)$ specified as the weights. To update the $\lambda$ estimates, we use logistic regression if $L$ is binary, and a multinomial log-linear model [48] if $J>2$, with $\xi_{ij}(t)$ used as weights. The EM process is repeated until convergence is reached. Once convergence has been attained, we use the Viterbi algorithm to predict the most likely sequence for $L_{t}$ given the estimated parameters and observed data.
@@ 4.2 Inference
In Appendix A we additionally consider a parametric bootstrap as was proposed by Visser et al. [53]). This approach, following the convergence of the HMM model, simulates a new multivariate time series of the observed variables using the obtained estimates and re-fits the HMM model for each bootstrap. We then obtain a percentile confidence interval from the bootstrap estimates. We denote this method as HMM-F-S-P for singular imputation of $Y_{t}$, HMM-F-M-P for the multiple imputation of $Y_{t}$, and HMM-F-ML-P for the multiple imputation of $Y_{t}$ and $L_{t}$, where the ‘P’ denotes parametric bootstrap. In Figure 6 we visually display the differences in the six proposed frequentist algorithms.
@@ 5 Simulation
We consider two simulation setting with $J=2$ or $J=3$ disease states, and $T=500$ follow-up points. We include two internal covariates $C_{t}=\{C_{1,t},C_{2,t}\}$ and two external covariates $W_{t}=\{W_{1,t},W_{2,t}\}$. We simulate $Y_{t},C_{1,t},C_{2,t}$, and $W_{1,t}$ as continuous with normally distributed errors. We let $W_{2,t}$ represent the indicator of it being a weekend night. For each setting we generate 200 simulations. We focus on a scenario which considers an $A_{t}$ as continuous with normal errors, however in Appendix A we explore the performance of the proposed methods in an additional scenario where $A_{t}$ is generated and modeled as an ordinal variable, with proportional odds of ascending to an additional state of $A_{t}$ assumed to be constant. We simulate $Y$ to be approximately 30% missing at random, and all other variables are fully observed. To each resulting multivariate time series, we first apply a naive linear model which assumes stationarity and does not identify nor account for modification by $L_{t}$. For this model, we obtain a 95% confidence interval using traditional asymptotic assumptions of normality. For the proposed methods, we complete 1000 bootstraps for each model type and calculate the empirical 95% confidence interval from the bootstraps. For the multiple imputation model, HMM-F-ML-B, we include ten multiple imputations. For the two proposed hidden Markov models, after convergence is reached we use the Viterbi algorithm to estimate the most likely sequence of $L_{t}$ and compare this to the true $L_{t}$ used to simulate the data. We then re-label the imputed states to maximize the possible prediction accuracy. This procedure addresses the potential that a model may flip the labels for the latent states, but still be accurate in prediction and estimating coefficients of interest.
@@
In Figure 2, we plot the results for the estimates of $\beta_{3,1}$, $\beta_{3,2}$, and $\beta_{3,3}$, the contemporaneous conditional average individual causal effect of $A_{t}=a_{t}+1$ versus $A_{t}=a_{t}$ on $Y_{t}$ given $L_{t}=1$, $L_{t}=2$, and $L_{t}=3$ respectively. The first row displays results from the scenario with two underlying latent states (only $\beta_{3,1}$ and $\beta_{3,2}$ estimated), while the second shows results with a multinomial $L_{t}$ with three underlying latent states. First most notably here, we see how the naive approach fails to estimate the effects of interest, as it does not account for modification by the latent state. We note that the estimates for both proposed models are very accurate and unbiased. In Figure 3, we show the confidence interval coverage, representing the percent of the 95% CI intervals which include the truth across simulations. As might be expected given the bias in the estimates for the naive approach, this method achieves 0% coverage. However, both proposed methods achieve acceptable coverage regardless of whether $L_{t}$ is binary or multinomial. While the estimates for the coefficient of $A$’s effect on $Y$ across the HMM models are consistently unbiased in the simulation, we note that several other estimates do tend to be slightly biased, likely due to multicollinearity.
@@ 6 Application
The Bipolar Longitudinal Study (BLS) is an ongoing mobile health study at McLean Hospital where 74 patients with schizophrenia or bipolar disorder are followed for up to five years. From the individual’s smartphone, their anonymized call and text logs, GPS movements, accelerometer behavior, and daily symptom, emotion, and behavior survey responses are recorded via the Beiwe application [54, 10]. The patients additionally participate in approximately monthly in-person visits where a range of assessments are conducted. Here, we consider a several distinct exposures of interest which are captured using both active ecological momentary assessment (EMA) data, and passive data to show how the proposed approach can allow for the integration of different data modalities.
Here we first consider a female participant with a bipolar spectrum disorder and a follow-up of 708 days. Our outcome of interest ($Y$) is overnight sleep duration, as identified from watch actigraph data via the deep phenotyping of sleep (DPSleep) algorithm [55]. We use daily total incoming and outgoing text count with key contacts to be the exposure, where key contacts were identified as individuals with whom the participant is in frequent reciprocated contact. To account for skewness, in the model we use a square root transformation on the exposure, meaning $A_{t}=\sqrt{\text{text count with key contacts}}$. We additionally include negative and positive mood in the model as covariates, $C_{1,t},C_{2,t}$, respectively. Negative and positive mood are calculated as an aggregate score from 0 to 28 of negative and positive items on the daily survey. Lastly, we use the local daily max environmental temperature as $W_{1,t}$, and whether the night is a weekday or weekend as $W_{2,t}$, covariates that are not affected by the participant. We apply our proposed modeling methods to the individual’s data, using the same specifications as in the simulation. We specify the model to include two underlying latent states, as clinically we understand this patient only exhibits symptoms related to depressed and euthymic states. We note that while $A_{t}$ and $W_{t}$ are fully observed, 23% of $C_{t}$ is missing, representing non-response for the daily EMA data. As the model currently is only designed to allow missing data in $Y$, we impute missing values in $C$ using Kalman smoothing imputation prior to fitting the model. Overnight sleep duration, $Y_{t}$ is missing 20% of the days across follow-up, and we leave this missingness to be handled by the models.
@@
In Figure 8 we plot the observed time series and model results for a male individual, Participant B, with bipolar spectrum disorder and a follow-up of 984 days. Here, while we include the same outcome and covariates as for Participant A, we define the exposure of interest using GPS signal. The exposure $A_{t}$ specifically denotes the percentage of time spent away from home on a given day. Notably this individual has a much higher rate of missing data, with 68% of EMA responses unobserved. Unfortunately, likely due to the high rate of missingness and low variabilitiy in the observed data, we see that the models are unable to capture clear trends for the latent disease state. While the frequentist model with single implementation (HMM-F-S-B) has rapid fluctuations between the two identified latent states, the model with multiple imputation (HMM-F-ML-B) has some more stable sections between periods with rapid fluctuations, but does not appear to capture meaningful disease states. Overall, this participant demonstrates a setting where the proposed methods are incapable of identifying a clinically relevant disease state, due to the high rate of missing data and lack of variability in the observed data.

### v2

@@ 4 Model implementation
We consider a frequentist hidden Markov model approach (HMM-F) to estimate the parameters defined in Section 3. In Appendix A, we explore an equivalent model fit within a Bayesian framework using a Markov Chain Monte Carlo sampling procedure. For the frequentist model, we implement an expectation-maximization (EM) procedure based on the Baum-Welch algorithm. In the expectation step, forward probabilities $\alpha_{i}(t)$ and backward probabilities $\beta_{i}(t)$ are calculated recursively using $a_{ij,t-1}$ and $b_{j,Y_{t},A_{t},C_{t}}(y_{t},a_{t},c_{t})$ [46, 28] such that
@@
To update the $\beta$, $\psi$, and $\phi$ estimates, we use the appropriate generalized linear regression for $Y_{t}$, $A_{t}$, and $C_{t}$ as the outcome with $\gamma_{i}(t)$ specified as the weights. To update the $\lambda$ estimates, we use logistic regression if $L$ is binary and a multinomial log-linear model [48] if $J>2$, with $\xi_{ij}(t)$ used as weights. The EM process is repeated until convergence is reached. Once convergence has been attained, we use the Viterbi algorithm to predict the most likely sequence for $L_{t}$ given the estimated parameters and observed data.
@@ 4.2 Inference
In Appendix A we additionally consider a parametric bootstrap as was proposed by Visser et al. [53]. This approach, following the convergence of the HMM model, simulates a new multivariate time series of the observed variables using the obtained estimates and re-fits the HMM model for each bootstrap. We then obtain a percentile confidence interval from the bootstrap estimates. We denote this method as HMM-F-S-P for singular imputation of $Y_{t}$, HMM-F-M-P for the multiple imputation of $Y_{t}$, and HMM-F-ML-P for the multiple imputation of $Y_{t}$ and $L_{t}$, where the ‘P’ denotes parametric bootstrap. In Figure 6 we visually display the differences in the six proposed frequentist algorithms.
@@ 5 Simulation
We consider two simulation setting with $J=2$ or $J=3$ disease states, and $T=500$ follow-up points. We include two internal covariates $C_{t}=\{C_{1,t},C_{2,t}\}$ and two external covariates $W_{t}=\{W_{1,t},W_{2,t}\}$. We simulate $Y_{t},C_{1,t},C_{2,t}$, and $W_{1,t}$ as continuous with normally distributed errors. We let $W_{2,t}$ represent the indicator of it being a weekend night. For each setting we generate 200 simulations. We focus on a scenario which considers an $A_{t}$ as continuous with normal errors, however in Appendix A we explore the performance of the proposed methods in an additional scenario where $A_{t}$ is generated and modeled as an ordinal variable, with proportional odds of ascending to an additional state of $A_{t}$ assumed to be constant. We simulate $Y$ to be approximately 30% missing at random, and all other variables are fully observed. To each resulting multivariate time series, we first apply a naive linear model which assumes stationarity and does not identify nor account for modification by $L_{t}$. For this model, we obtain a 95% confidence interval using traditional asymptotic assumptions of normality. For the proposed methods, we complete 1000 bootstraps for each model type and calculate the empirical 95% confidence interval from the bootstraps. For the multiple imputation model, HMM-F-ML-B, we include ten multiple imputations. For the two proposed hidden Markov models, after convergence is reached we use the Viterbi algorithm to estimate the most likely sequence of $L_{t}$ and compare this to the true $L_{t}$ used to simulate the data. We then re-label the imputed states to maximize the possible prediction accuracy. This procedure addresses the potential that a model may switch the labels for the latent states, but still be accurate in prediction and estimating coefficients of interest.
@@
In Figure 2, we plot the results for the estimates of $\beta_{3,1}$, $\beta_{3,2}$, and $\beta_{3,3}$, the contemporaneous conditional average individual causal effect of $A_{t}=a_{t}+1$ versus $A_{t}=a_{t}$ on $Y_{t}$ given $L_{t}=1$, $L_{t}=2$, and $L_{t}=3$, respectively. The first row displays results from the scenario with two underlying latent states (only $\beta_{3,1}$ and $\beta_{3,2}$ estimated), while the second shows results with a multinomial $L_{t}$ with three underlying latent states. Most notably here, we see how the naive approach fails to estimate the effects of interest, as it does not account for modification by the latent state. We note that the estimates for both proposed models are very accurate and unbiased. In Figure 3, we show the confidence interval coverage, representing the percent of the 95% CI intervals which include the truth across simulations. As might be expected given the bias in the estimates for the naive approach, this method achieves 0% coverage. However, both proposed methods achieve acceptable coverage regardless of whether $L_{t}$ is binary or multinomial. While the estimates for the coefficient of $A$’s effect on $Y$ across the HMM models are consistently unbiased in the simulation, we note that several other estimates do tend to be slightly biased, likely due to multicollinearity.
@@ 6 Application
The Bipolar Longitudinal Study (BLS) is an ongoing mobile health study at McLean Hospital where 74 patients with schizophrenia or bipolar disorder are followed for up to five years. From the individual’s smartphone, their anonymized call and text logs, GPS movements, accelerometer behavior, and daily symptom, emotion, and behavior survey responses are recorded via the Beiwe application [54]. The patients additionally participate in approximately monthly in-person visits where a range of assessments are conducted. Here, we consider a several distinct exposures of interest which are captured using both active ecological momentary assessment (EMA) data, and passive data to show how the proposed approach can allow for the integration of different data modalities.
Here we first consider a female participant with a bipolar spectrum disorder and a follow-up of 708 days. Our outcome of interest ($Y$) is overnight sleep duration, as identified from watch actigraph data via the deep phenotyping of sleep (DPSleep) algorithm [55]. We use daily total incoming and outgoing text counts with key contacts as the exposure, where key contacts were identified as individuals with whom the participant is in frequent reciprocated contact. To account for skewness, in the model we use a square root transformation on the exposure, meaning $A_{t}=\sqrt{\text{text count with key contacts}}$. We additionally include negative and positive mood in the model as covariates, $C_{1,t},C_{2,t}$, respectively. Negative and positive mood are calculated as an aggregate score from 0 to 28 of negative and positive items on the daily survey. Lastly, we use the local daily max environmental temperature as $W_{1,t}$, and whether the night is a weekday or weekend as $W_{2,t}$, covariates that are not affected by the participant. We apply our proposed modeling methods to the individual’s data, using the same specifications as in the simulation. We specify the model to include two underlying latent states, as clinically we understand this patient only exhibits symptoms related to depressed and euthymic states. We note that while $A_{t}$ and $W_{t}$ are fully observed, 23% of $C_{t}$ is missing, representing non-response for the daily EMA data. As the model currently is only designed to allow missing data in $Y$, we impute missing values in $C$ using Kalman smoothing imputation prior to fitting the model. Overnight sleep duration, $Y_{t}$ is missing 20% of the days across follow-up, and we leave this missingness to be handled by the models.
@@
In Figure 8 we plot the observed time series and model results for a male individual, Participant B, with bipolar spectrum disorder and a follow-up of 984 days. Here, while we include the same outcome and covariates as for Participant A, we define the exposure of interest using GPS. The exposure $A_{t}$ specifically denotes the percentage of time spent away from home on a given day, which is calculated using the DPLocate pipeline [56]. Notably this individual has a much higher rate of missing data, with 68% of EMA responses unobserved. Unfortunately, likely due to the high rate of missingness and low variability in the observed data, we see that the models are unable to capture clear trends for the latent disease state. While the frequentist model with single implementation (HMM-F-S-B) has rapid fluctuations between the two identified latent states, the model with multiple imputation (HMM-F-ML-B) has some more stable sections between periods with rapid fluctuations, but does not appear to capture meaningful disease states. Overall, this participant demonstrates a setting where the proposed methods are incapable of identifying a clinically relevant disease state, due to the high rate of missing data and lack of variability in the observed data.
@@ 7 Discussion
Acknoledgements
We thank Dost Ongur for his valuable insights and thoughtful contributions.